OpenStudy (anonymous):

What is the value of the dependent variable if the value of the independent variable is –4? f(x)=-3x2+5x

OpenStudy (anonymous):

I'm assuming this means x=-4 because of the definition of IV/DV f(x)=-3(-4)^2+5(-4) f(x)=-48-20 f(x)=-68